# Near-inertial echoes of ageostrophic instability in submesoscale filaments

### v1

@@ 4 Results
Figure 2: Evolution of the mean state for the simulation with $Ri_{\text{min}}=0.1$. Three snapshots are shown at $ft/2\pi=\{0.8,1.5,2.2\}$. The heatmap shows $\overline{v}$. Black contours are mean isopyncals $\overline{b}$. Orange (purple) contours are (anti-)clockwise streamlines.
@@ 4.1 General evolution
The reference filament state does not satisfy the boundary conditions and there is an initial adjustment phase that creates an Ekman layer at the surface. For the first 2–3 inertial periods, the filament remains approximately uniform in the down-front direction. Towards the end of our simulations, and especially visible at the surface, the isopycnals begin to meander due to baroclinic instability. The domain is not large enough to support many baroclinic modes; following Stone (1966), the fastest-growing baroclinic mode wavelength $\lambda$ can be estimated for an Eady model with parameters comparable to the centre of a jet
$\lambda=\frac{2\pi v_{0}}{f}\sqrt{\frac{1+Ri_{\min{}}}{5/2}}\approx 5,$ (8)
@@
which is half the downfront length of our domain. We ran a verification simulation with a longer down-front domain that resolved many wavelengths of this instability (not shown), and found that the baroclinic instability develops too slowly to be significant in the time period of interest in this paper. The rest of this discussion is concerned with the evolution of the filament before the meandering and shortly after the adjustment at the surface, which can be understood from the perspective of the down-front mean state.
Figure 2 outlines the evolution of the mean state during the first inertial periods. Shortly after the filament is allowed to evolve freely, a strong secondary circulation on the scale of the fronts emerges (figure 2a). The secondary circulation acts to restratify the fronts. At the surface, the circulation increases front strength: horizontal convergence pushes isopycnals together immediately under the surface (figure 2b), before the trend reverses somewhat and fronts weaken (figure 2c). The maximum horizontal buoyancy gradient does not occur in the centre of the filament, but at approximately $x=\pm 0.4$. The evolution is contained in the region with negative potential vorticity and the fronts behave somewhat independently, rather than as a filament.
This happens within less than an inertial period and therefore, the frontal strengthening and arrest mechanisms do not appear to be balanced at some maximum front strength. This is very suggestive of a damped oscillation within the filament.
@@ 4.2 Initial evolution is a fast instability
In this section, we show that the dynamics in the first inertial period contains a low-$q$ instability of the vertical shear. The form of the instability consists of thin rolls aligned with isopyncals in vertical ($x$-$z$) slices that extract energy from the mean vertical shear via a negative VSP. However, there is significant down-front variation of the flow.
Figure 3: Terms extracting energy from the mean state for each of the simulations integrated over the mixed layer.
Figure 4: At $t/2\pi=0.25$, comparison of an $x$-$z$ slice at $y=0$ of a) vertical velocity $w$ and b) vertical shear production in the left side of the filament for the simulation with $Ri_{\text{min}}=0.1$. The red, dashed line in the top left panel locates $z=-0.05$, namely, the depth of the $x$-$y$ slice of vertical velocity displayed in c). The horizontal slice of $w$ is coarse-grained with a Gaussian filter with a half-width of one grid cell for clarity. A video timeseries of this plot appears in the supplementary material as Movie 1.
Figure 5: As in figure 4 but at $t/2\pi=0.5$. Note the difference in vertical velocity and shear production scales compared to figure 4.
@@
The instantaneous vertical velocity fields show the form of the instability. Figures 4 and 5 show the VSP and horizontal ($x$-$y$) and vertical ($x$-$z$) slices of the vertical velocity $w$ at $t/2\pi=0.25$ and $0.5$. In the vertical slices, the velocity field consists of thin cells, aligned with isopycnals. A horizontal slice through the emerging cells (figure 4c) reveals that they are not aligned along the down-front direction, as would be expected of SI. The cells appear rotated away from the symmetric axis. Later (figure 5) cells emerge that appear to align with the filament and resemble an expected SI circulation. Videos of the vertical velocity and VSP as shown in figures 4 and 5 are presented in the supplementary material.
Once the instability saturates and gives way to turbulence, there is a large production of turbulent kinetic energy via negative VSP. The saturated regions in figure 4a,b above $z=-0.05$ are associated with a larger VSP signature. The maximum in VSP moves down over time, and figure 5c shows the horizontal state as the coherent cells are destroyed by secondary instabilities and the turbulence begins to isotropise in the $x$-$y$ plane. The extraction of energy is fast, and can be seen in figure 3 to grow, peak and subside within the first inertial period. The consequence is that the filament is now out of geostrophic balance, as shown in the following section.
@@ 4.3 The initial instability echoes throughout the front as an inertial oscillation
Following the onset of instability, the evolution of the filament is not balanced. Indeed, the process in the previous section happens within less than an inertial period, and in the resulting evolution the front is not stabilised at some minimum width. We highlight the consequences of the short-lived turbulent energy production production using the linearised Sawyer-Eliassen equation for the mean secondary circulation streamfunction $\psi$ forced by by-products of turbulent vertical fluxes as we measure them in our simulations. That is, we use
$\nabla^{2}\psi_{tt}+\underbrace{f^{2}\psi_{zz}-J(fv_{0},\psi_{z})+J(b_{0},\psi_{x})}_{L\psi}=\underbrace{f\overline{w^{\prime}v^{\prime}}_{zz}-\overline{w^{\prime}b^{\prime}}_{xz}}_{G},$ (9)
@@
$(\overline{u},0,\overline{v})=(-\psi_{z},0,\psi_{x}),\quad\text{and}\quad J(a,b)=a_{z}b_{x}-b_{z}a_{x}.$ ($)10𝑎𝑏$
@@
$\frac{\text{d}^{2}C_{\text{IML}}}{\text{d}t^{2}}=\int_{-0.9H}^{-0.1H}\text{d}z\int^{0}_{-L_{x}/2}\text{d}x\Big{[}f^{2}\psi_{zz}-J(fv_{0},\psi_{z})+J(b_{0},\psi_{x})-f\overline{w^{\prime}v^{\prime}}_{zz}+\overline{w^{\prime}b^{\prime}}_{xz}\Big{]}.$ (11)
The frontal circulation appears dominated by decaying near-inertial oscillations. The peak in VSP and thus $\overline{w^{\prime}v^{\prime}}$ is large, and occurs in less than one inertial period, before significant change in the mean state; this suggests that the subsequent behaviour will be unbalanced and lead to oscillations. A comparison of the terms in equation (11) is shown for each of the simulations in figure 6. The circulation shows clear decaying oscillatory behaviour with period $2\pi$, with linear terms dominating for small $Ri_{\min{}}$. Frontal arrest is not necessarily due to horizontal diffusion, but rather the restoring phase of the oscillation. The strength of these oscillations is dependent on the Richardson number of the flow: a smaller $Ri_{\min{}}$ produces a faster and stronger extraction of kinetic energy and in turn leads to higher amplitude oscillations.
Figure 6: The circulation around the left half of the interior mixed layer as in equation (11). The black line is the actual evolution of the circulation. The blue line is the contribution of the linear terms, the orange line is the contribution of the by-products of turbulent flux terms, and the green line is their sum. $L\psi$ and $G$ are defined as in equation (9)
@@ 5 Discussion and conclusions
We investigated the evolution of low-Richardson number mixed layer filaments with numerical simulations. A low-$q$ instability quickly extracts energy from the vertical shear. This puts the filament in an unbalanced state, and results in vertically sheared inertial oscillations. The oscillations are well-described by the linear terms in the Sawyer-Eliassen equation for the down-filament streamfunction, following a forcing due to the vertical turbulent flux of down-front momentum.
Vertical mixing caused by fast-acting instability in a front is capable of producing secondary circulations superficially similar to those produced when a front is forced by a strain flow or turbulence. Rather than solely restratifying the mixed layer, the instability at least initially strengthens surface buoyancy gradients.
We believe that this initial circulation is at least partially responsible for the filamentogenesis observed by Sullivan & McWilliams (2018). The authors integrate up to slightly past maximum front strength, so would not see inertial oscillations resulting from an instability-induced imbalance. Here we run smaller, yet still substantial, simulations of a similar setup and observe a similar behaviour up to that point in time, without any forcing by surface cooling or winds.
Wienkers et al. (2021) observe a similar instability-to-inertial-oscillations pathway that begins with a strong symmetric instability in 2.5D direct numerical simulations of finite-width, zero Richardson number frontal zones (i.e., constant lateral buoyancy gradients in domains that are periodic in the cross-front direction). The behaviour of our filament with $Ri_{\min{}}=0$ is remarkably similar. For instance, the leftmost plot of figure 3 shows reoccuring creation of turbulent kinetic energy via negative VSP associated with the steepest isopycnals at each phase of the inertial oscillation, much like figure 4 in the above reference. Our simulations are three-dimensional and reveal that, while it shares features of SI, the instabilities that eventually lead to inertial oscillations have significant down-front variation, at least initially.
@@
Submesoscale structures in a realistic, time-varying ocean can show inertial oscillations associated with the diurnal variation of vertical diffusivity as shown by Dauhajre & McWilliams (2018). We demonstrate here that the changes in effective diffusivity need not come from changes in surface forcing, but can arise without forcing, from turbulent kinetic energy production due to submesoscale instabilities acting on a front.
@@
Frontal features created in the prototypical sense, via a confluent strain flow, cross a range of scales before they reach the regime of the filament presented here. We believe that this work indicates there is understanding to be gained by spinning up fronts in-situ for idealised numerical studies of turbulence, rather than beginning with what we think a submesoscale front ought to look like. In particular, the transition from $Ro\ll 1$ to $Ro\sim 1$ frontogenesis as these scales remain poorly resolved by global ocean models.
@@
Submesoscale fronts and filaments are the sites of various instabilities, the consequences on the global ocean and best parametrisation of which remains an open question. Here we have shown that instability alone is capable of strengthening surface buoyancy gradients in a mixed layer filament via a multi-stage process in which the initial instability acts quickly, puts the fronts out of balance and the resulting adjustment begins with a convergent secondary circulation on the scale of the fronts that strengthen them near the surface. For our experiments, frontal arrest is not necessarily due to turbulence and can be described simply as the restoring phase of an inertial oscillation. The instability in question shares characteristics with SI, in agreement with previous studies, but the initial emergence has significant down-front variation, before symmetric modes appear later. We also comment on the consequences for studies of submesoscale flows, and suggest that the most fruitful way to investigate how these structures evolve in the real ocean, from a computational perspective, may be to grow, rather than impose, analogues.
@@ Appendix A Parametrisation of the reference state
$h(x)=H\gamma\left(\frac{x}{\ell}\right)\quad\text{with}\quad\gamma(s)=-1+\frac{\delta}{2}\left[\text{erf}\left(s+\frac{1}{2\alpha}\right)-\text{erf}\left(s-\frac{1}{2\alpha}\right)\right].$ ($)13𝑎𝑏$
@@
$Ro_{\text{min}}=\frac{(N^{2}-N_{0}^{2})\beta^{2}}{f^{2}\alpha^{2}}\max_{s}\Big{(}\gamma\gamma_{s}\Big{)}\quad\text{and}\quad Ri_{\text{min}}=\frac{f^{2}\alpha^{2}N^{2}}{\beta^{2}(N^{2}-N_{0}^{2})^{2}}\min_{s}\Big{(}\gamma_{s}^{-2}\Big{)}.$ ($)14𝑎𝑏$

### v2

@@ 3.1 General evolution
For the first 2–3 inertial periods ($\sim 54~{}\text{hr}$ at mid-latitudes), the filament remains approximately uniform in the down-front direction. Towards the end of our simulations ($ft/2\pi=5.4$, not shown), and especially visible at the surface, the isopycnals begin to meander due to baroclinic instability. The domain is not large enough to support many baroclinic modes; following Stone (1966), the fastest-growing baroclinic mode wavelength $\lambda$ can be estimated for an Eady model with parameters comparable to the centre of a jet,
$\lambda=\frac{2\pi v_{0}}{f}\sqrt{\frac{1+Ri_{\min{}}}{5/2}}\approx 5L,\quad$ (7)
@@
which is half the downfront length of our domain. The small domain limits the baroclinic growth rate. Nevertheless, we ran a simulation with a longer down-front domain that resolved around eight wavelengths of this instability (not shown), in which the meandering indeed becomes noticeable sooner, at around $ft/2\pi\approx 2$. Nevertheless, we verified that even then, the baroclinic instability does not alter the conclusions we present in this study. The rest of this discussion is concerned with the evolution of the filament before the meandering, which can be understood from the perspective of the down-front mean state.
Figure 3: Evolution of the mean state for the simulation with $Ri_{\text{min}}=0.0$. Three snapshots are shown at $ft/2\pi=\{0.8,1.5,2.2\}$. The heatmap shows $\overline{v}$. Solid black contours are mean isopyncals $\overline{b}$. Dashed orange (purple) contours are (anti-)clockwise streamlines of the down-front mean flow.
Figure 4: a) A Hovmöller plot of the horizontal buoyancy gradient of the down-front mean state for the simulation with $Ri_{\text{min}}=0.0$, averaged over the top 10% of the mixed layer. Black contours are contours of the buoyancy averaged over the same region. Red, dashed lines indicate the times at which the snapshots in figure 3 are taken. b) The down-front mean sub-grid dissipation of kinetic energy $\overline{\varepsilon}=\overline{\nu\nabla\boldsymbol{u}:\nabla\boldsymbol{u}}$, as in a).
The vertically sheared jet does not satisfy the no-flux boundary condition at $z=0$ for the down-front velocity and there is an onset of restratifying across-front Ekman flow; however by $ft/2\pi=0.7$ this is overshadowed by a strong secondary circulation throughout the mixed layer, which is the focus of this section.
Figure 3 outlines the evolution of the down-front mean state during the first inertial periods. Shortly after the filament is allowed to evolve freely, a strong secondary circulation on the scale of the fronts emerges (figure 3a). The secondary circulation acts to restratify the fronts. At the surface, the circulation increases front strength: horizontal convergence pushes isopycnals together immediately under the surface. Afterwards, the secondary circulation changes sign (figure 3b). Hovmöller plots of the down-front-averaged horizontal buoyancy gradient and sub-grid dissipation, averaged over the down-front direction and the top 10% of the mixed layer are shown in figure 4. Near-surface isopycnals are drawn together, and horizontal buoyancy gradients reach a maximum strength at $ft/2\pi\approx 1.3$. This is associated with the maximum level of sub-grid dissipation. Afterwards, the down-front-averaged near-surface buoyancy gradients weaken before somewhat strengthening again, reaching a second maximum around one inertial period after the initial maximum. A sub-grid dissipation signal also recurs at this time ($ft/2\pi\approx 2.1$). The fronts approach a final state (disregarding the later meandering due to baroclinic instability) with surface buoyancy gradients that are stronger than those in the reference state, though weaker than the maxima achieved earlier.
@@ 3.2 Initial evolution is a fast instability
In this section, we show that the dynamics in the first inertial period contains an instability that extracts energy from the the vertical sheared velocity.
Figure 5: Terms extracting energy from the mean state for each of the simulations integrated over the mixed layer. Each term is normalised by the maximum VSP in the simulation with $Ri_{\min{}}=0.0$. Marked points in c) indicate the times at which the snapshots of vertical velocity and contribution to VSP in figures 6 and 7 are taken.
Figure 6: At $t/2\pi=0.30$, comparison of an $x$-$z$ slice at $y=0$ of a) vertical velocity $w$ and b) contribution to vertical shear production in the left side of the filament for the simulation with $Ri_{\text{min}}=0.1$. The red, dashed line in the top left panel locates $z=-50\text{m}$, the depth of the $x$-$y$ slice of vertical velocity displayed in c). The horizontal slice of $w$ is coarse-grained with a Gaussian filter with a half-width of one grid cell for clarity. A video timeseries of this plot appears in the supplementary material as Movie 1.
Figure 7: As in figure 6 but at $t/2\pi=0.46$, the time at which VSP is maximum (see 5). Note the difference in vertical velocity and shear production scales compared to figure 6.
@@
Figures 6 and 7 show the VSP and horizontal ($x$-$y$) and vertical ($x$-$z$) slices of the vertical velocity $w$ at $ft/2\pi=0.30$ and $0.46$ respectively for one of the fronts that comprises the filament. At the earlier time, the contribution of the turbulence to vertical shear production is greater near the surface, and the horizontal slice shows some organisation within the front, though there is significant down-front variation.
Once the instability saturates and gives way to turbulence, there is a large production of turbulent kinetic energy via VSP. The saturated regions in figure 6a,b above $z=-50\text{ m}$ are associated with a larger VSP signature. The maximum in VSP moves down over time, indicating that the instability grows in a non-normal fashion, and figure 7c shows the horizontal state as the somewhat coherent cells are destroyed by secondary instabilities and the turbulence begins to isotropise in the $x$-$y$ plane. The extraction of energy is fast, and can be seen in figure 5 to grow, peak and subside within the first inertial period. The consequence is that the filament is now out of geostrophic balance, as shown in the following section.
A video of the vertical velocity and mean contribution to VSP as shown in figures 6 and 7 is presented in the supplementary material. Appendix B presents changes in behaviour as the large initial noise (figure 2) is lowered. Doing so recovers the symmetric normal modes as expected of the low-Ri initial conditions. While qualitatively distinct, the effect of the normal-mode SI on the filament is largely the same as the high-noise case, insofar as this study is concerned. That is, even though normal-mode SI grows somewhat more slowly than the non-normal instability we just described, it remains fast enough to cause a rapid imbalance of the jet, followed by a filament-scale inertial oscillation, similar to the one we are about to describe.
@@ 3.3 The initial instability echoes throughout the front as an inertial oscillation
Following the onset of instability, the evolution of the filament is not balanced. Indeed, the process in the previous section happens within less than an inertial period, and in the resulting evolution the front does not stabilise at some minimum width. We highlight the consequences of the short-lived turbulent energy production using the linearised Sawyer-Eliassen equation for the mean secondary circulation streamfunction $\psi$ forced by by-products of turbulent vertical fluxes as we measure them in our simulations. That is, we use
$\nabla^{2}{\psi}_{\!,{tt}}+\underbrace{f^{2}{\psi}_{\!,{zz}}-J(fv_{0},{\psi}_{\!,{z}})+J(b_{0},{\psi}_{\!,{x}})}_{L}=\underbrace{f{\overline{w^{\prime}v^{\prime}}}_{\!,{zz}}-{\overline{w^{\prime}b^{\prime}}}_{\!,{xz}}}_{G},$ (8)
@@
$(\overline{u},0,\overline{v})=(-{\psi}_{\!,{z}},0,{\psi}_{\!,{x}}),\quad\text{and}\quad J(a,b)={a}_{\!,{z}}{b}_{\!,{x}}-{b}_{\!,{z}}{a}_{\!,{x}}.$ ($)9𝑎𝑏$
@@
$\frac{\text{d}^{2}C_{\text{IML}}}{\text{d}t^{2}}=\int_{-0.9H}^{-0.1H}\text{d}z\int^{0}_{-L_{x}/2}\text{d}x\Big{[}f^{2}{\psi}_{\!,{zz}}-J(fv_{0},{\psi}_{\!,{z}})+J(b_{0},{\psi}_{\!,{x}})-f{\overline{w^{\prime}v^{\prime}}}_{\!,{zz}}+{\overline{w^{\prime}b^{\prime}}}_{\!,{xz}}\Big{]}.$ (10)
The frontal circulation appears dominated by decaying near-inertial oscillations. The peak in $G$ is large, and occurs in less than one inertial period, before significant change in the mean state; this indicates that the subsequent behaviour will be unbalanced and lead to oscillations. To confirm, a comparison of the terms in equation (10) is shown for each of the simulations in figure 8. The circulation shows clear decaying oscillatory behaviour with period $2\pi/f$, with linear terms dominating for small $Ri_{\min{}}$. The strength of these oscillations is dependent on the Richardson number of the flow: a smaller $Ri_{\min{}}$ produces a faster and stronger extraction of kinetic energy and in turn leads to higher amplitude oscillations.
Figure 8: The circulation around the left half of the interior mixed layer as in equation (10). The black line is the actual evolution of the circulation; i.e. the left side of (10). The blue line is the contribution of the linear terms, the orange line is the contribution of the by-products of turbulent flux terms, and the green line is their sum. $L$ and $G$ are defined as in equation (8). All timeseries are smoothed with a Gaussian kernel with standard deviation $0.1/f$.
The secondary circulation is correlated with the surface front strength. The maximum front strength in figure 4a emerges at $ft/2\pi\approx 1$, at the same time that isopycnals begin to move apart. This coincides with the moment $\frac{\text{d}^{2}C}{\text{d}t^{2}}=0$, and the circulation changes sign (figure 8a).
@@ 4 Discussion and conclusions
We investigated the evolution of low-Richardson number mixed layer filaments with numerical simulations. Energy is quickly extracted, via vertical shear production, from the down-front velocity. This puts the filament in an unbalanced state, and results in vertically sheared inertial oscillations that strengthen horizontal buoyancy gradients near the surface, while weakening them near the base of the mixed layer. The oscillations are well-described by the linear terms in the Sawyer-Eliassen equation for the down-front streamfunction, following a forcing induced by the vertical turbulent flux of down-front momentum.
The specific nature of the instability is outside the scope of this paper, but will no doubt be of interest to readers. In short, we believe it to be a non-normal, transient response towards SI that is sensitive to the initial noise, rather than normal-mode SI. While we leave a full demonstration to future studies, we find a comparison with reduced-noise simulations particularly illuminating. We present these results in appendix B. The initial noise illustrated in figure 2 has a large magnitude, compared to the mean flow; and its vertical profile is responsible for the apparent downward propagation of the VSP contribution from figures 6 to 7. At low levels of initial noise, the unstable modes that emerge are down-front symmetric (figure 9) before being destroyed by secondary instabilities (not shown). Perturbations emerge at the same rate throughout the mixed layer in a normal-mode fashion, rather than growing fastest at the surface. The low-noise behaviour is similar to the uniform-gradient Eady state studied by Wienkers et al. (2021). The authors observe a similar instability-to-inertial-oscillations pathway that begins with a strong symmetric instability in 2.5D direct numerical simulations of finite-width, zero Richardson number frontal zones (i.e., constant lateral buoyancy gradients in domains that are periodic in the cross-front direction).
Arobone & Sarkar (2015), Zemskova et al. (2020), and Kimura (2024) investigate the initial and long-time behaviour of a symmetrically unstable flow. The initial and normal growth of perturbations in such a flow are different. In particular, the perturbations are initially strongly asymmetric and not aligned with isopycnals, before becoming nearly symmetric for large times (Arobone & Sarkar, 2015). The instability at the start of the simulations in this paper is such an initial response of the filament associated with SI and, due to the large initial noise, the perturbations do not have time to grow into normal-mode SI before being isotropised by secondary instabilities acting on the growing modes, leading to the rapid energy extraction from the mean state via the VSP.
@@
Submesoscale structures in a realistic, time-varying ocean can show inertial oscillations associated with the diurnal variation of vertical diffusivity as shown by Dauhajre & McWilliams (2018). We demonstrate here that the changes in effective diffusivity need not come from changes in surface forcing, but can arise without forcing, from turbulent kinetic energy production due to submesoscale instabilities acting on a front. The resulting inertial oscillations contribute to surface frontal strengthening, but originate from unstable initial conditions, rather than a general feature of mixed layer frontogenesis in the real ocean, where instabilities, fronts, and turbulent mixing co-exist and evolve together. With similarly unstable initial conditions, and the additional presence of a persistent turbulence-driven secondary circulation, we suspect that the inertial oscillations will be superimposed on a monotonic, turbulent thermal wind frontogenesis.
@@
Frontal features created in the prototypical sense, via a confluent strain flow, cross a range of scales before they reach the regime of the filament presented here. We believe that this work indicates there is understanding to be gained by spinning up fronts in-situ for idealised numerical studies of turbulence, rather than beginning with what we think a submesoscale front ought to look like. In particular, the transition from $Ro\ll 1$ to $Ro\sim 1$ frontogenesis as these scales remain poorly resolved by global ocean models. Of particular interest here, Thomas (2012) showed that the behaviour of near-inertial oscillations is significantly altered in regions of frontogenesis driven by strain. Near-inertial waves are captured by growing fronts, which quickly rotate the wavevector to align with isopycnals, before kinetic energy extraction via the frontogenetic secondary circulation damps them completely. Therefore, we believe that in large-eddy simulations combining vertical convection and mixed layer strain-driven frontogenesis across scales, near-inertial oscillations of the nature described in this paper will not be present.
@@
Submesoscale fronts and filaments are the sites of various instabilities, the consequences on the global ocean and best parametrisation of which remains an open question. Here we have shown that instability alone is capable of strengthening surface buoyancy gradients in a mixed layer filament via a multi-stage process in which the initial instability, which is likely a non-normal, transient response to initial noise, acts quickly and puts the fronts out of balance. The resulting adjustment begins with a convergent secondary circulation on the scale of the fronts that strengthens them near the surface. For our experiments, arrest of surface convergence is not exclusively due to turbulence and is at least in part the restoring phase of the inertial oscillation. We also comment on the consequences for studies of submesoscale flows, and suggest that the most fruitful way to investigate how these structures evolve in the real ocean, from a computational perspective, may require spinning up — rather than imposing — initial conditions, e.g. starting with a larger scale mixed layer front and applying an advective forcing, to avoid energising unwanted dynamics due to instability or otherwise.
@@ Appendix A Parametrisation of the reference state
$h(x)=H\gamma\left(\frac{x}{\ell}\right)\quad\text{with}\quad\gamma(s)=-1+\frac{\delta}{2}\left[\text{erf}\left(s+\frac{1}{2\alpha}\right)-\text{erf}\left(s-\frac{1}{2\alpha}\right)\right].$ ($)12𝑎𝑏$
@@
$Ro_{\text{min}}=\frac{(N^{2}-N_{0}^{2})\beta^{2}}{f^{2}\alpha^{2}}\max_{s}\Big{(}\gamma\gamma_{s}\Big{)}\quad\text{and}\quad Ri_{\text{min}}=\frac{f^{2}\alpha^{2}N^{2}}{\beta^{2}(N^{2}-N_{0}^{2})^{2}}\min_{s}\Big{(}\gamma_{s}^{-2}\Big{)}.$ ($)13𝑎𝑏$
@@
Appendix B Transient behaviour and the impact of initial noise
In the main body of this paper, we outline the process leading to inertial oscillations in an unstable front. In this additional section, we will attempt to connect the topic of this study with symmetric instability literature. We find that the emergence and propagation of the initial disturbance throughout the filament is dependent on the amplitude and form of the fluctuations in the initial conditions. This is beyond the scope of the main body in this paper, which is concerned with the consequences of unstable initial conditions for frontogenesis studies, but we include this secondary set of simulations and analyses to provide a more thorough exploration of the generation of inertial oscillations, as well as to demonstrate that the conclusions in the main body of the text hold for rather different circumstances.
Due to the choice of pre-initial conditions, the initialisation described in section 2 produces a state with a turbulent kinetic energy profile that is approximately linear in $z$ (from zero at the base of the mixed layer to a maximum near the thermocline). This serves as an analogue for the TKE of a state driven by surface cooling. This profile is the cause of the downward propagation of the VSP from figure 6 to 7. Fluctuations simply start larger at the surface. Altering the noise profile to be constant in depth in the mixed layer, for example, causes the VSP to emerge approximately at the same time throughout the unstable region (not shown).
Here we present a set of three simulations, each with different values for the maximum (surface) amplitude of the pre-initial noise. The simulation in the main body of this paper has maximum pre-initial noise amplitude a factor $10^{-2}$ of the maximum down-front velocity in the reference state. In this section we repeat the analysis in the main body of the paper, but for simulations pre-initialised with noise of relative magnitude $10^{-4}$, $10^{-6}$, $10^{-8}$.
For the simulation with smallest pre-initial noise amplitude, the form of the vertical velocity at early times is shown in figure 9. Characteristic SI modes are recovered. These modes are somewhat tilted with respect to the isopycnals, expected for non-hydrostatic SI (Wienkers et al., 2021).
Figure 9: a, b) vertical slices of vertical velocity of the simulation with smallest pre-initial noise amplitude ($10^{-8}$) at a) $ft/2\pi=0.8$ and b) $ft/2\pi=1.0$ the time of maximum $\text{VSP}_{0}$ (see figure 10). The red, dashed line in the top right panel locates $z=-50\text{ m}$, the depth of the $x$-$y$ slice of vertical velocity displayed in c). c) a horizontal slice of vertical velocity at $z=-50\text{ m}$ at the time of maximum $\text{VSP}_{0}$.
The down-front mean defined in the earlier sections does not remove the large across-front wavenumber symmetric modes and as such is not appropriate nor illustrative to describe a coarse-grained state. With this in mind, and for this section only, we define the vertical and lateral shear productions, as well as the vertical buoyancy flux, using the departure of the simulation state from the reference filament state, namely,
$\text{LSP}_{0}=-\int_{\mathcal{V}}\text{d}V\;\ \frac{\partial\boldsymbol{u}_{0}}{\partial x}\boldsymbol{\cdot}u(\boldsymbol{u}-\boldsymbol{u}_{0}),\quad\text{VSP}_{0}=-\int_{\mathcal{V}}\text{d}V\;\ \frac{\partial\boldsymbol{u}_{0}}{\partial z}\boldsymbol{\cdot}w(\boldsymbol{u}-\boldsymbol{u}_{0}),$ (14)
and
$\text{BFLUX}_{0}=\int_{\mathcal{V}}\text{d}V\;w(b-b_{0}).$ (15)
We continue with these alternate definitions, noting that the perturbations from the reference state will include the down-front-mean ageostrophic circulation and may only be appropriate at early times. Nevertheless, figure 5 was reproduced for the first set of simulations with the above definitions (not shown) and bears only quantitative differences.
Figure 10 shows the shear production and buoyancy flux terms for a set of simulations with decreasing amplitude of initial noise. In addition, it includes the VSP as defined in equation (6). The $\text{VSP}_{0}$ is the primary source of energy for the fluctuations, and behaviour across all cases up to the extremum in $\text{VSP}_{0}$ is largely identical; though, for the largest two noise amplitudes ($10^{-2}$, in figure 5 and $10^{-4}$), the $\text{VSP}_{0}$ growth and peak occurs earlier. As expected, VSP associated with the perturbations vanishes as they become more symmetric at lower noise levels.
Figure 10: As in figure 5 but for a series of simulations with decreasing amplitudes of initial noise relative to the maximum velocity, and using the TKE production definitions in equations (14) and (15). From left to right, the simulations are pre-initialised with noise of relative magnitudes $10^{-4}$, $10^{-6}$, and $10^{-8}$. Terms are shown relative to the same scale in figure 5. Additionally, and for comparison, the VSP calculated according to equation (6) is included as a dashed green line. Marked points in c) indicate the times at which the snapshots of vertical velocity in figure 9 are taken.
Following the TKE production by $\text{VSP}_{0}$, inertial oscillations are also observed in the secondary circulations, as seen in figure 11. Hovmöller plots in figure 12 show that the near-surface isopycnals follow the inertial oscillations, with a large front strength and enhanced dissipation just after the filament reaches a minimum width ($ft/2\pi\approx 1.7$). The generation of inertial oscillations and subsequent surface convergence and frontal strengthening is agnostic to the method of extraction of kinetic energy, as long as it is rapid enough.
Figure 11: As in figure 8 but for a series of simulations with decreasing amplitudes of initial noise relative to the maximum velocity. From left to right, the simulations are pre-initialised with noise of relative magnitude $10^{-4}$, $10^{-6}$, and $10^{-8}$. All timeseries are smoothed with a Gaussian kernel with standard deviation $0.3/f$.
Figure 12: a) A Hovmöller plot of the horizontal buoyancy gradient of the down-front mean state for the simulation with smallest pre-initial noise amplitude ($10^{-8}$), averaged over the top 10% of the mixed layer. Black contours are contours of the buoyancy averaged over the same region. b) The down-front mean sub-grid dissipation of kinetic energy $\overline{\varepsilon}=\overline{\nu\nabla\boldsymbol{u}:\nabla\boldsymbol{u}}$, as in a).